Business and finance. In planning for a new item, a manufacturer assumes that the number of items produced x and the cost in dollars C of producing these items are related by a linear equation. Projections are that 100 items will cost $10,000 to produce and that 300 items will cost $22,000 to produce. Find the equation that relates C and x.
Answer by stanbon(48510)

You have two points: (100,10000) and (300,22000)
The equation of the relationship will be C=mx+b
Find "m": m = [22000-10000]/[300-100]= 12000/200= 60
You now have x=100 when C=10000, and m=60
Put those into the form C=mx+b to find "b".
10000=60(100)+b
10000=6000+b
b=4000
The equation you want is C=60x+4000
Cheers,
Stan H.